OpenStudy (anonymous):

What is an example of a 2-step equation with no Solutions?

OpenStudy (anonymous):

11 + 3x – 7 = 6x + 5 – 3x 11 +6x = 6x + 5 11 = 5 6x + 5 = 6x 5 = 0

OpenStudy (anonymous):

either one should do. Probably the second one

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem! (anymore)

OpenStudy (anonymous):

:3 Nice